OpenStudy (anonymous):

Difference of Squares in Trigonometric Functions \[\frac{ 3tanx }{ 1+\tan^2 }=3sinxcosx\]

OpenStudy (anonymous):

Are you trying to prove that the LHS = RHS ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

what are you stuck with? Do you understand how to approach this problem?

OpenStudy (anonymous):

I mean, I understand that you need to factor in order to solve it but I don't know where to start

OpenStudy (anonymous):

\[(3tanx)/(1+\tan^2x) = 3sinxcosx\] \[Since, tanx = sinx/cosx\] We can do the following: \[3sinx/cosx/(1+ \sin^2x/\cos^2x)\]

OpenStudy (anonymous):

then find a common denominator which is cos^x

OpenStudy (anonymous):

\[3sinx/cosx * \cos^2x / \cos^2x+\sin^2x\] Then we notice that cos^2 + sinx^2 = 1 <- which is the Pythagorean identity.

OpenStudy (anonymous):

then after more algebraic simplification we get 3sinxcosx

OpenStudy (anonymous):

Ahh thank you!

OpenStudy (anonymous):

thus both sides match and the identity holds.

OpenStudy (anonymous):

Your welcome.